OpenStudy (here_to_help15):

A Triangle has side lengths of 6,8,10.Is it a right triangle? Explain

OpenStudy (luffingsails):

If it is a right triangle, then the pythagorean theorem will be true. Try it and see.

OpenStudy (here_to_help15):

No plagarism please

OpenStudy (luffingsails):

is 6^2 + 8^2 = 10^2 ?

OpenStudy (here_to_help15):

Ok so how would i break this down

OpenStudy (here_to_help15):

?

OpenStudy (luffingsails):

Just as I wrote... try the Pythagorean theorem using the 3 sides of your triangle. If the equation is false, then the triangle is NOT right.

OpenStudy (here_to_help15):

thats where im confused lol

OpenStudy (luffingsails):

6^2 + 8^2 = 10^2? 36 + 64 = 100? The answer is yes. So, it is a right triangle!

OpenStudy (here_to_help15):

I may be great at everything else just no Geo V.V

OpenStudy (luffingsails):

That's okay. I completely understand. With me it's probabilities. lol

OpenStudy (here_to_help15):

lol that also o.o

OpenStudy (luffingsails):

One tool we have to confirm right triangles (outside of trig) is the Pythagorean theorem. Easy to apply.

OpenStudy (here_to_help15):

Ok bro :D

OpenStudy (luffingsails):

Cheers. Make sense?

OpenStudy (here_to_help15):

Yesh

OpenStudy (luffingsails):

Cool. good morning.

OpenStudy (here_to_help15):

you are a he correct... lol

OpenStudy (here_to_help15):

@luffingsails